OpenStudy (fuzzydragon):

If Bob throws a water balloon out of his dorm room window, which is 39 feet off the ground, and the initial velocity is 96 ft/sec, how long until the water balloon hits the ground? Use h=vt-16t^2

jimthompson5910 (jim_thompson5910):

in this case, h = 39 and v = 96. Plug those values in and solve for t

OpenStudy (fuzzydragon):

That wouldn't work though because I am trying to find the time at height 0 (when it hits the ground)

jimthompson5910 (jim_thompson5910):

then maybe h = 0 will work

OpenStudy (fuzzydragon):

h=0 and still the velocity of 96

jimthompson5910 (jim_thompson5910):

but the issue is that we don't account for the 39 ft

jimthompson5910 (jim_thompson5910):

hmm

jimthompson5910 (jim_thompson5910):

maybe the equation should be h(t) = vt-16t^2+h

OpenStudy (fuzzydragon):

The equation in the question is the one my teacher gave

OpenStudy (fuzzydragon):

So I don't think that I am allowed to use a different one

jimthompson5910 (jim_thompson5910):

alright

jimthompson5910 (jim_thompson5910):

then use h = 0 and v = 96 and see if it works

OpenStudy (fuzzydragon):

\[0=96t-16t^2\] \[16t^2-96t=0\] \[t^2-6t=0\] \[(t+2)(t-3)=0\] \[t=-2 t=3\]

OpenStudy (fuzzydragon):

there should be a space between t=-2 and t=3

jimthompson5910 (jim_thompson5910):

t^2-6t=0 does NOT turn into (t+2)(t-3)=0

jimthompson5910 (jim_thompson5910):

t^2-6t=0 turns into t(t-6) = 0

OpenStudy (fuzzydragon):

Yeah sorry

OpenStudy (fuzzydragon):

so t=0 and t=6

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

although, i'm not sure about the equation you're given, so idk 100%

OpenStudy (fuzzydragon):

Except my teacher had said that it would not factor evenly and we'd have to use the quadratic formula

OpenStudy (fuzzydragon):

So I am quite stuck.

jimthompson5910 (jim_thompson5910):

so maybe there's a typo somewhere?

OpenStudy (fuzzydragon):

Our notes packet says \[h=vt-4.9t\] for meters and \[h=vt-16t^2\] for feet

jimthompson5910 (jim_thompson5910):

I have a feeling it's something like h=vt-16t^2 39=96t-16t^2 which would lead you to have to use the quadratic formula, but again not 100% sure

OpenStudy (fuzzydragon):

That should be \[h=vt-4.9t^2\]

jimthompson5910 (jim_thompson5910):

but I'm more used to h(t) = -16t^2 + vt + h

jimthompson5910 (jim_thompson5910):

so h(t) = -16t^2 + vt + h h(t) = -16t^2 + 96t + 39 then you plug in h(t) = 0 and solve for t

OpenStudy (fuzzydragon):

I am looking in my notes and the example equation is "A ball is thrown directly upward from ground level with an initial velocity of 80 ft/sec. When will the ball return to the ground? How high will the ball go?" h=80t-16t^2 0=16t(t-5) t=0, t=5 h=80(2.5)-16(2.5)^2 h=200-100 100 feet

jimthompson5910 (jim_thompson5910):

notice how this one is thrown from the ground level (height = 0)

jimthompson5910 (jim_thompson5910):

but for this one says "Bob throws a water balloon out of his dorm room window, which is 39 feet off the ground" the starting height is 39 ft

OpenStudy (fuzzydragon):

So you do think that I should use 39=96t-16t^2

jimthompson5910 (jim_thompson5910):

more like 0 = -16t^2 + 96t + 39

jimthompson5910 (jim_thompson5910):

0 = -16t^2 + vt only applies when you start on the ground

jimthompson5910 (jim_thompson5910):

that +39 bumps everything up 39 so you start up 39 ft

OpenStudy (fuzzydragon):

So instead of equaling 39 (which would end up subtracting it), we add 39?

jimthompson5910 (jim_thompson5910):

yeah solve 0 = -16t^2 + 96t + 39 for t

jimthompson5910 (jim_thompson5910):

and the quadratic formula will have to be used

OpenStudy (fuzzydragon):

Okay, thank you so much!

jimthompson5910 (jim_thompson5910):

np